OpenStudy (lgbasallote):

a tank contains 100 gal. of water. brine enters the tank at the rate of 3 gpm. the mixture, thoroughly stirred, leaves the tank at the rate of 2 gpm. if the concentration of the brine at the end of 20 minutes is to be 2 lb/gal, what should be the concentration of the brine entering the tank? Ans. 4.75 lb/gal

OpenStudy (lgbasallote):

how to set this up?

OpenStudy (lgbasallote):

is it still ambiguous @Mimi_x3 ?

OpenStudy (lalaly):

let S be salt in tank at time t F1 flow rate with salt concentration M1 \[\frac{ds}{dt}=F_1M_1-F_2M_2\]\[\frac{ds}{dt}=3M_1-2(\frac{s}{100+(3-2)t})\]\[\frac{ds}{dt}=3M_1-\frac{2s}{100+t}\]\[\frac{ds}{dt}+\frac{2s}{100+t}=3M_1\]now solve using integrating factor method\[I=e^{\int\limits{\frac{2}{100+t}dt}}=e^{2\ln(100+t)}=(100+t)^2\]multiply by DE\[(100+t)^2\frac{ds}{dt}+2s(100+t)=3M_1(100+t)^2\]notice that RHS is product rule\[\frac{d}{dt}((100+t)^2 \times s)=3M_1(100+t)^2\]integrate both sides\[(100+t)^2s=M_1(100+t)^3+C\]\[s=\frac{M_1(100+t)^3+C}{(100+t)^2}\]at t=0 --> S=0 \[C=-1\times 10^6M_1\]so\[s=\frac{M_1(100+t)^3+1 \times 10^6M_1}{(100+t)^2}\]at t=20 M1=2 lb/gal\[2=\frac{s}{100+20}\]\[s=240\] \[240=\frac{M_1(100+20)^3+1 \times 10^6M_1}{(100+20)^2}\]solving for M1 \[M_1=4.75 lb/gal\]

OpenStudy (lalaly):

should be -1x10^6M1

OpenStudy (lalaly):

@lgbasallote

OpenStudy (lgbasallote):

i got lost after "multiply with DE" what did you do there?

OpenStudy (lalaly):

The right hand side looks like the product rule\[(uv)'=u'v+uv'\]

OpenStudy (lalaly):

do u get it?

OpenStudy (lgbasallote):

hmm yup seems you did the long way of linear DE got me confused..im used to shortcut... t -> 0 then s ->0 does this always apply?

OpenStudy (lalaly):

yeah initial value

OpenStudy (lgbasallote):

so if s is not given then at t = 0 it's always 0? that means it starts with fresh water? no solution inside the tank or something?

OpenStudy (lalaly):

yeah the liquid has not enterd the tank yet so there is no solution in the tank

OpenStudy (lgbasallote):

okay you said c = -1 x 10^6 M1 but in the next step you did \[s = \frac{M_1 (100+t)^3 + 1 \times 10^6 M_1}{(100+t)^2}\] is this a typo?

OpenStudy (lgbasallote):

okay i follow some of the others except that...and how you got 4.75?

OpenStudy (lalaly):

yeah typo thats why i reposted it in the second post

OpenStudy (lgbasallote):

oh lol didnt see that

OpenStudy (lgbasallote):

so how did you get 4.75? can you show the substitution?

OpenStudy (phi):

*Mt(t^2+300t+30000) +c

OpenStudy (lgbasallote):

hmm that's the same as \[\large Q(100 + t)^2 = Mt(100 + t)^3 + C\]

OpenStudy (lgbasallote):

right?

OpenStudy (phi):

are you asking if Mt(t^2+300t+30000)=? Mt(100+t)^3 or (t^2+300t+30000)=? (100+t)^3

OpenStudy (lgbasallote):

im asking if what you said is the same as what i said

OpenStudy (lgbasallote):

because i cannot see the division of the equal sign...idk which corres[ponds to which

OpenStudy (lgbasallote):

aww why

OpenStudy (phi):

start with \[ \frac{d}{dt}((100+t)^2 \times s)=3M_1(100+t)^2 \] or \[ d ((100+t)^2 s)= 3M(100+t)^2 dt \]

OpenStudy (lgbasallote):

\[\int 3M_1 (100 + t)^2 dt \implies M_1 (100+t)^3\] right?

OpenStudy (phi):

yes, it's right. I was looking for why I was not getting M1=4.75 from the last equation. Looking at it, it looks like a sign got flipped on the C value. so, with that typo fixed, all is ok

OpenStudy (lgbasallote):

so how to get 4.75?

OpenStudy (phi):

replace s with 240, t=20 (change +M*10^6 to -M*10^6) , factor out the M...

OpenStudy (lgbasallote):

hmm...mind writing it?

OpenStudy (phi):

\[ s = \frac{M_1 (100+t)^3 - 1 \times 10^6 M_1}{(100+t)^2} \] \[ s=\frac{(100+t)^3 - 10^6}{(100+t)^2 }M_1 \] with s= 240 and t=20 \[ 240= \frac{120^3-10^6}{120^2} M_1 \]

OpenStudy (lgbasallote):

ohhh i see... i was multiplying instead of dividing lol

OpenStudy (lgbasallote):

thanks for clarifying @phi :D

OpenStudy (lgbasallote):

@phi one more question...what was the question asking for?

OpenStudy (lgbasallote):

was it the salt concentration at t = 0? or t = 20?

OpenStudy (phi):

the salt concentration of the entering solution at t=20

OpenStudy (lgbasallote):

so what is 2 lb/gal??

OpenStudy (phi):

2 lb/gal is the concentration of salt in the tank at t=20 It starts at t=0 with pure water, and brine enters (4.75 lbs/gal) at 3 gal/min and at the same time, the solution leaves the tank at 2 gal/min

OpenStudy (lgbasallote):

ohhh 2 lb/gal is *in* the tank 4.75 is the one going inside the tank?

OpenStudy (phi):

4.75 is the brine going into the tank. At some point, when the tank reaches equilibrium, we would expect 4.75 lbs salt per gal in the tank, leaving the tank and entering the tank

OpenStudy (lgbasallote):

so is 4.75 constant? or it only applies at t=20?

OpenStudy (phi):

the 4.75 lbs/gal is the concentration of brine going into the tank. It is constant, and is the amount that leads to 2 lbs/gal of salt in the tank after exactly 20 minutes.

OpenStudy (lgbasallote):

ohhhh

OpenStudy (lgbasallote):

last one... is 2 lb/gal concentration? or amt of salt? im confused with that part

OpenStudy (phi):

I am being sloppy. 2 lbs of salt per gallon of brine (salt+water) is a concentration, from which we can deduce the amount of salt in the tank.

OpenStudy (lgbasallote):

ohhh i see thanks!!!

OpenStudy (lgbasallote):

ohhh..it's making sense to me now slowly...i thought concentration only comes in..it's also possible that there is already concentration in the tank already right?